# Joint Probability Distribution of mRNA and Protein Molecules in a Stochastic Gene Expression Model

### v2

@@ 3.2 Reconstruct Probability Mass Function
Under certain condition, all moments can uniquely determine the underlying probability distribution [CourseProbabilityTheory2000]. Particularly, the binomial moments may determine the probability mass function through a compact identity. According to (7), stationary probability mass function $\mathbb{P}(m,n)$ can be reconstructed from binomial moments according to
@@
We note that, in practice, computation based on (9) suffers from severe floating-point error if not implemented properly. This is because for large $p$ and $q$, multiplying the relatively small binomial moment with the extremely large combinatorial coefficients can introduce significant floating-point errors. This is essentially the same numerical issue encountered and analyzed in [Exactlysolvablemodels2020J.Chem.Phys., lu2025stochastickineticsmrnamolecules]. In the accompanying codes, we take the approach in [lu2025stochastickineticsmrnamolecules], carrying out computations in the logarithmic domain and exponentiating the final results.
We analyze an example using both the analytical results (8), SSA, and FSP; the results are shown in Figure 2, Figure 3, and Figure 4, respectively. The mean squared error between the joint probability mass function in Figure 2 and Figure 3 is $1.9208\times 10^{-8}$; and between the joint probability mass function in Figure 2 and Figure 4 is $1.7685\times 10^{-12}$. See the accompanying Jupyter Notebook for detail.
@@ 3.3 Low-Order Moments
Using the hierarchal approach of calculating the binomial moments described in the above section, we can readily obtain explicit expressions of low-order binomial moments. For simplicity, we now stop tracking the state of gene by considering only the coarse-grained binomial moments $B_{p,q}:=\bm{\pi}^{\top}\mathcal{B}_{p,q}\bm{1}$ and the probability mass function $P(m,n):=\bm{\pi}^{\top}\mathbb{P}(m,n)\bm{1}$ for $p,q,m,n\in\mathbb{N}$, similar to [lu2025stochastickineticsmrnamolecules]. $\bm{1}\in\mathbb{R}^{N\times 1}$ denotes the all-ones vector. $\bm{\pi}\in\mathbb{R}^{N\times 1}$ is the invariant distribution of the underlying Markov chain $(S(t))_{t\geq 0}$ characterized by $D$. Assume $D$ is irreducible. Then $\bm{\pi}$ is the unique vector satisfying $\bm{\pi}^{\top}D=\bm{0}_{1\times N}$ and $\bm{\pi}^{\top}\bm{1}=1$. Explicit expressions of $\mathcal{B}_{p,q}$, $B_{p,q}$ with $L\leq 2$ and the detailed derivation can be found in the Supplementary Material. For comparison with binomial moments derived from models under burst approximation in section 4, we particularly present here the first two binomial moments of protein copy number.
@@ 4 Accuracy of Burst Approximation
Burst approximation is commonly used to simplify the complete gene expression model (1). Intuitively, burst approximation aims to capture the experimental phenomena that proteins are often produced in bursts. In one burst, proteins are synthesized at a relatively high rate over a short interval, followed by a long period of silence. One important quantitative relation underlying such burst phenomena is that, in general, mRNAs have much shorter lifetimes than proteins. In the notations of (1), this corresponds to $u\gg\delta$. However, it has been argued that $u\gg\delta$ is far from sufficient to generate genuine translation bursts, and bust approximation is better understood as a conceptual mathematical technique [Modelsstochasticgene2005Phys.LifeRev.].
Under burst approximation, mRNA molecules can be eliminated formally from the system, and only the one-dimensional probability distribution of protein molecules needs to be revealed. As a result, stochastic models under burst approximation are typically easier to analyze. For models with one active gene state and multiple inactive states, a compact analytical expression for the probability mass function can be derived in terms of the generalized hypergeometric function [Exactdistributionsstochastic2022Phys.Rev.E, Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. By contrast, this is generally impossible for the complete gene expression model (2) according to the analytical expression of the generating function (See the Supplementary Material). General models with multiple gene states are also studied under burst approximation, and a recurrence relation for binomial moments can also be obtained using one-dimensional binomial moment method [Exactdistributionsstochastic2022Phys.Rev.E]. Additionally, non-Markovian models (models that are not continuous-time Markov chains) under burst approximation exist and can be studied using queueing theory [Intrinsicnoisestochastic2011Phys.Rev.Lett.]. We note that, according to the equivalence between queueing system and gene expression model established in [Solvingstochasticgeneexpression2024Biophys.J.], models under burst approximation can be interpreted as queueing systems with batch arrivals [Analysisinfiniteserverqueue2002QueueingSyst.]. This perspective may explain the significant convenience introduced by the burst approximation.
In general, stochastic gene expression models under burst approximation possess independent theoretical value. However, in this article, we focus on the validity of burst approximation when replacing the original reaction system (1) with an alternative. Therefore, the general setting is not introduced here. A systematic analysis of general stochastic gene expression models under burst approximation will be reported in a forthcoming paper [lu2025stochastickineticsproteinmolecules].
@@
$\rightarrowfill@\displaystyle{}\mathrm{{Protein}}{}\mathrel{\mathop{\makebox[0.0pt]{\rightarrowfill@\displaystyle}}\limits^{\mkern 5.0mu{}\mathrm{{}}\mkern 9.0mu}}{}{}\mathrm{\text{$\emptyset$}}$
In (11), the species ${S}_{i}\;(1\leq i\leq N)$, Protein, and the parameters $a_{i,j}\;(i\neq j,\;\;1\leq i,j\leq N)$, $\delta$ carry the same meaning as in (1). Additionally, under the notations in (1), $b_{i,j}^{[r]}\;(1\leq i,j\leq N,\;\;r=1,2,\cdots)$ are given by the following relation.
@@
The relation (12) can be interpreted as follows. Once transcribed, an mRNA molecule is subject to two competing reaction pathways, namely, hydrolysis and translation. More specifically, this can be seen as the competing binding of decay complexes that promote hydrolysis, and recruitment of initiation factors that engage the ribosome for translation. Since the probability of initiating translation rather than hydrolysis is $v/(u+v)$, the number of protein molecules produced from a single mRNA molecule follows a geometric distribution with parameter $u/(u+v)$. Hence, (12) readily follows. We note that the geometrically distributed burst size is consistent with experimental observations [Stochasticproteinexpression2006Naturea].
With the notations introduced before, the first-order and second-order binomial moments of protein copy number in (11) are
@@
The detailed derivation can be found in [lu2025stochastickineticsproteinmolecules].
Compared with (10), it is worth noting that the first-order binomial moment, equivalently the expectation of the protein copy number, remains exact under burst approximation. By contrast, the second-order binomial moment is explicitly altered when burst approximation is applied. In particular, an upper bound can be derived for the difference between the binomial moments obtained from the two models, using techniques from functional analysis and an analogue to Theorem 4.1.2 in [MatrixComputations2013]. Specifically, we have
@@
Note that the infinity norm of a matrix, denoted by $\lVert\cdot\rVert_{\infty}$, is the maximum absolute row sum of this matrix. The proof can be found in the Supplementary Material. According to (14), the difference converges to zero at the rate of $1/u^{2}$ if $u\rightarrow\infty$ while the other parameters are fixed. In general, $u/\delta\gg 1$ does not guarantee the validity of the buret approximation. In the left panel of Figure 5, both the upper bound given by (14) and the gap between variances converges to zero as $u$ grows and all other parameters remain the same. However, according to the right panel of Figure 5, although $u/\delta\gg 1$, the gap between variances of the protein copy number in complete gene expression models and the corresponding surrogate models can be arbitrarily large, indicating that the burst approximation can be inaccurate in some cases.
Figure 5: Gap between Variance and its Upper Bound: In the left panel, all the parameters expect for $u$ are the same as those in Figure 2 and are fixed. While $u$ varies in $[1,20]$, we compute the discrepancy between the variances of the protein copy number according to (10) and Equation 13, and also evaluate its upper bound given by (14). In the right panel, $u=10$, $v=2$, and $\delta=1$ are fixed, while $D_{0}$ and $D_{1}$ take the form $a_{i,j}=0\;(i\neq j)$ and $b_{i,j}=i$. For the order of the model $1\leq N\leq 100$, we compute the discrepancy between the variances of the protein copy number according to (10) and Equation 13, and also evaluate its upper bound given by (14). Note that both illustrations are plotted in $\ln(\cdot)$ scale, and that the discrepancy between variances equals two times the discrepancy between second-order binomial moments.
@@ 5 Discussion and Conclusion
In this paper, we establish an effective approach to analyzing stochastic gene expression models without resorting to burst approximation. Analytical expressions of binomial moments of mRNA and protein counts can be obtained up to any order at steady state in a hierarchical manner. Numerical computation of binomial moments is both fast and accurate. Subsequently, joint probability mass function can be reconstructed. In particular, based on analytical expressions of low-order cumulants, we can rigorously evaluate the validity of the burst approximation for general gene expression models.
@@
We first note that asymptotic behavior of binomial moments and probability mass function needs further analysis. In [lu2025stochastickineticsmrnamolecules], we derive elegant upper bounds for binomial moments and probability mass function of mRNA copy number, which are $B_{p,0},\;p\in\mathbb{N}$ and $\sum_{n=0}^{\infty}P(m,n),\;m\in\mathbb{N}$ in this article, respectively. However, deriving concise upper bounds for $B_{p,q},\;p,q\in\mathbb{N}$ or $P(m,n),\;m,n\in\mathbb{N}$ appears substantially more involved. Note that asymptotic analysis is crucial to proving the convergence of (9) and designing an appropriate truncation strategy. A rigorous treatment is left to future work.
@@
In the end, we examine the gene-state topology from a multiscale modeling perspective. Given the flexibility of our general model (1), it becomes essential to predefine the number of states $N$ and the transition structure, namely $D_{0}$ and $D_{1}$. Otherwise, the model may yield highly pathological results and nonphysical predictions. Fortunately, from a multiscale modeling perspective, this task may be accomplished by molecular dynamics (MD) simulations [StatisticalMechanicsTheory2023]. At the atomistic level, dynamics of macromolecules, such as DNA and proteins, can be accurately captured using MD simulations. Physically, each gene state corresponds to a metastable configuration of the underlying macromolecule, namely DNA in this problem. Thus, the Markov chain of gene state is a coarse-grained representation of the underlying dynamics. Designing the gene-state topology essentially amounts to constructing a Markov state model (MSM) from MD trajectories, which is a standard method for analyzing MD data and understanding conformational transitions [Identificationslowmolecular2013J.Chem.Phys., MetastabilityMarkovState2013]. Constructing a MSM typically involves two steps. The first step is to project the high-dimensional MD data, namely the coordinates of all atoms in a macromolecule, onto a relatively lower-dimensional latent space. The above dimensionality reduction relies on specifying collective variables (CVs), which map the complete configurations to low-dimensional representations but preserve non-trivial dynamical behaviors of the molecule. Next, one clusters the data in CV space and uses cluster centroids as discrete states of a Markov chain, whose transition rates are estimated at a given lag time. Nevertheless, constructing CVs remains a challenge for real-world macromolecules, although machine learning techniques have been incorporated to develop many useful methods [Collectivevariablediscovery2022RSCAdv., unifiedframeworkmachine2023J.Chem.Phys.]. Therefore, further work is needed to develop a physically grounded, bottom-up stochastic gene expression model.
@@ Appendix A Matrix-form Chemical Master Equation and Generating Function Method
In this section, we show that the chemical master equation (2) can be equivalently reformulated as the partial differential equation system (4) using standard generating function method.
@@ Appendix D Proof of (14)
In the penultimate inequality, we applied Hölder’s inequality. In the last inequality, we used an analogue to Theorem 4.1.2 in [MatrixComputations2013], stating that
@@
Since $D$ is the generator of a continuous-time Markov chain, for any $\lambda>0$, the matrix $\lambda\bm{I}_{N}-D$ is strictly row diagonally dominant, with each row summing to $\lambda$. Therefore $\lVert\lambda\bm{I}_{N}-D\rVert_{\infty}\leq 1/\lambda$.
@@ Appendix E Analytical Expression of Generating Function
where $H(s;z,w;t):=D_{0}+h(s;z,w,t)D_{1}\in\mathbb{R}^{N\times N}$ and $h(s;z,w;t)$ is a real-valued function whose expression is given below.
@@ Appendix F Hierarchical solver for $\mathcal{B}_{p,q}$ by Layers
3:function SolveRight(M,YM,\,Y) $\triangleright$ Return $X$ such that XM=YX\,M=Y (Use LU factorization [MatrixComputations2013])
@@
12:  $Z_{L}\leftarrow\textsc{SolveRight}(M_{n},\ -\,Y_{L-1}D_{1})$

### v3

@@ 3.2 Reconstruct Probability Mass Function
Under certain conditions, all moments can uniquely determine the underlying probability distribution [7]. Particularly, the binomial moments may determine the probability mass function through a compact identity. According to (7), stationary probability mass function $\mathbb{P}(m,n)$ can be reconstructed from binomial moments according to [1, 2, 38]
@@
We note that, in practice, computation based on (9) suffers from severe floating-point error if not implemented properly. This is because for large $p$ and $q$, multiplying relatively small binomial moment by extremely large combinatorial coefficients can introduce significant floating-point errors. This is essentially the same numerical issue encountered and analyzed in [13, 22]. In the accompanying codes, we take the approach in [22], carrying out computations in the logarithmic domain and exponentiating the final results.
We analyze an example using both the analytical results (8), SSA, and FSP; the results are shown in Figure 2, Figure 3, and Figure 4, respectively. The mean squared error between the joint probability mass function in Figure 2 and Figure 3 is $1.9208\times 10^{-8}$; and between the joint probability mass function in Figure 2 and Figure 4 is $1.7685\times 10^{-12}$. See the accompanying Jupyter Notebook for details.
@@ 3.3 Low-Order Moments
Using the hierarchical approach of calculating the binomial moments described in the above section, we can readily obtain explicit expressions of low-order binomial moments. For simplicity, we now stop tracking the state of the gene by considering only the coarse-grained binomial moments $B_{p,q}:=\bm{\pi}^{\top}\mathcal{B}_{p,q}\bm{1}$ and the probability mass function $P(m,n):=\bm{\pi}^{\top}\mathbb{P}(m,n)\bm{1}$ for $p,q,m,n\in\mathbb{N}$, similar to [22]. $\bm{1}\in\mathbb{R}^{N\times 1}$ denotes the all-ones vector. $\bm{\pi}\in\mathbb{R}^{N\times 1}$ is the invariant distribution of the underlying Markov chain $(S(t))_{t\geq 0}$ characterized by $D$. Assume $D$ is irreducible. Then $\bm{\pi}$ is the unique vector satisfying $\bm{\pi}^{\top}D=\bm{0}_{1\times N}$ and $\bm{\pi}^{\top}\bm{1}=1$. Explicit expressions of $\mathcal{B}_{p,q}$, $B_{p,q}$ with $L\leq 2$ and the detailed derivation can be found in the Supplementary Material. For comparison with binomial moments derived from models under burst approximation in section 4, we particularly present here the first two binomial moments of protein copy number.
@@ 4 Accuracy of Burst Approximation
The burst approximation is commonly used to simplify the complete gene expression model (1). Intuitively, the burst approximation aims to capture the experimental phenomenon that proteins are often produced in bursts. In one burst, proteins are synthesized at a relatively high rate over a short interval, followed by a long period of silence. One important quantitative relation underlying such burst phenomena is that, in general, mRNAs have much shorter lifetimes than proteins. In the notations of (1), this corresponds to $u\gg\delta$. However, it has been argued that $u\gg\delta$ is far from sufficient to generate genuine translation bursts, and the bust approximation is better understood as a conceptual mathematical technique [27].
Under burst approximation, mRNA molecules can be eliminated formally from the system, and only the one-dimensional probability distribution of the protein molecules needs to be determined. As a result, stochastic models under burst approximation are typically easier to analyze. For models with one active gene state and multiple inactive states, a compact analytical expression for the probability mass function can be derived in terms of the generalized hypergeometric function [6, 33]. By contrast, this is generally impossible for the complete gene expression model (2) according to the analytical expression of the generating function (See the Supplementary Material). General models with multiple gene states are also studied under burst approximation, and a recurrence relation for binomial moments can also be obtained using one-dimensional binomial moment method [6]. Additionally, non-Markovian models (models that are not continuous-time Markov chains) under burst approximation exist and can be studied using queueing theory [18]. We note that, according to the equivalence between queueing systems and gene expression models established in [35], models under burst approximation can be interpreted as queueing systems with batch arrivals [24]. This perspective may explain the significant convenience introduced by the burst approximation.
In general, stochastic gene expression models under burst approximation possess independent theoretical value. However, in this article, we focus on the validity of burst approximation when replacing the original reaction system (1) with an alternative. Therefore, the general setting is not introduced here. A systematic analysis of general stochastic gene expression models under burst approximation will be reported in a forthcoming paper [23].
@@
$\rightarrowfill@\displaystyle{}\mathrm{{Protein}}{}\mathrel{\mathop{\makebox[20.00003pt]{\rightarrowfill@\displaystyle}}\limits^{\mkern 5.0mu{}\mathrm{{}}\mkern 9.0mu}}{}{}\mathrm{\text{$\emptyset$}}$
In (11), the species $\mathcal{S}_{i}\;(1\leq i\leq N)$, Protein, and the parameters $a_{i,j}\;(i\neq j,\;\;1\leq i,j\leq N)$, $\delta$ carry the same meaning as in (1). Additionally, under the notations in (1), $b_{i,j}^{[r]}\;(1\leq i,j\leq N,\;\;r=1,2,\cdots)$ are given by the following relation.
@@
The relation (12) can be interpreted as follows. Once transcribed, an mRNA molecule is subject to two competing reaction pathways, namely, hydrolysis and translation. More specifically, this can be seen as the competing binding of decay complexes that promote hydrolysis, and recruitment of initiation factors that engage the ribosome for translation. Since the probability of initiating translation rather than hydrolysis is $v/(u+v)$, the number of protein molecules produced from a single mRNA molecule follows a geometric distribution with parameter $u/(u+v)$. Hence, (12) readily follows. We note that the geometrically distributed burst size is consistent with experimental observations [5].
With the notation introduced above, the first-order and second-order binomial moments of protein copy number in (11) are
@@
The detailed derivation can be found in [23].
Compared with (10), it is worth noting that the first-order binomial moment, equivalently the expectation of the protein copy number, remains exact under the burst approximation. By contrast, the second-order binomial moment is explicitly altered when burst approximation is applied. In particular, an upper bound can be derived for the difference between the binomial moments obtained from the two models, using techniques from functional analysis and an analogue to Theorem 4.1.2 in [12]. Specifically, we have
@@
Note that the infinity norm of a matrix, denoted by $\lVert\cdot\rVert_{\infty}$, is the maximum absolute row sum of this matrix. The proof can be found in the Supplementary Material. According to (14), the difference converges to zero at the rate of $O(u^{-2})$ as $u\rightarrow\infty$ while the other parameters are fixed. In general, $u/\delta\gg 1$ does not guarantee the validity of the burst approximation. In the left panel of Figure 5, both the upper bound given by (14) and the gap between the variances converge to zero as $u$ grows and all other parameters remain the same. However, according to the right panel of Figure 5, although $u/\delta\gg 1$, the gap between variances of the protein copy number in complete gene expression models and the corresponding surrogate models can be arbitrarily large, indicating that the burst approximation can be inaccurate in some cases.
Figure 5: Gap between Variance and its Upper Bound: In the left panel, all the parameters except for $u$ are the same as those in Figure 2 and are fixed. While $u$ varies in $[1,20]$, we compute the discrepancy between the variances of the protein copy number according to (10) and (13), and also evaluate its upper bound given by (14). In the right panel, $u=10$, $v=2$, and $\delta=1$ are fixed, while $D_{0}$ and $D_{1}$ take the form $a_{i,j}=0\;(i\neq j)$ and $b_{i,j}=i$. For the order of the model $1\leq N\leq 100$, we compute the discrepancy between the variances of the protein copy number according to (10) and (13), and also evaluate its upper bound given by (14). Note that both illustrations are plotted in $\ln(\cdot)$ scale, and that the discrepancy between variances equals two times the discrepancy between second-order binomial moments.
@@ 5 Discussion and Conclusion
In this paper, we establish an effective approach to analyzing stochastic gene expression models without resorting to burst approximation. Analytical expressions of binomial moments of mRNA and protein counts can be obtained up to any order at steady state in a hierarchical manner. Numerical computation of binomial moments is both fast and accurate. Subsequently, the joint probability mass function can be reconstructed. In particular, based on analytical expressions of low-order cumulants, we can rigorously evaluate the validity of the burst approximation for general gene expression models.
@@
We first note that the asymptotic behavior of binomial moments and probability mass function needs further analysis. In [22], we derive elegant upper bounds for binomial moments and probability mass function of mRNA copy number, which are $B_{p,0},\;p\in\mathbb{N}$ and $\sum_{n=0}^{\infty}P(m,n),\;m\in\mathbb{N}$ in this article, respectively. However, deriving concise upper bounds for $B_{p,q},\;p,q\in\mathbb{N}$ or $P(m,n),\;m,n\in\mathbb{N}$ appears substantially more involved. Note that asymptotic analysis is crucial to proving the convergence of (9) and designing an appropriate truncation strategy. A rigorous treatment is left to future work.
@@
Finally, we examine the gene-state topology from a multiscale modeling perspective. Given the flexibility of our general model (1), it becomes essential to predefine the number of states $N$ and the transition structure, namely $D_{0}$ and $D_{1}$. Otherwise, the model may yield highly pathological results and nonphysical predictions. Fortunately, from a multiscale modeling perspective, this task may be accomplished by molecular dynamics (MD) simulations [36]. At the atomistic level, dynamics of macromolecules, such as DNA and proteins, can be accurately captured using MD simulations. Physically, each gene state corresponds to a metastable configuration of the underlying macromolecule, namely DNA in this problem. Thus, the Markov chain of gene state is a coarse-grained representation of the underlying dynamics. Designing the gene-state topology essentially amounts to constructing a Markov state model (MSM) from MD trajectories, which is a standard method for analyzing MD data and understanding conformational transitions [28, 32]. Constructing a MSM typically involves two steps. The first step is to project the high-dimensional MD data, namely the coordinates of all atoms in a macromolecule, onto a relatively lower-dimensional latent space. The above dimensionality reduction relies on specifying collective variables (CVs), which map the complete configurations to low-dimensional representations but preserve non-trivial dynamical behaviors of the molecule. Next, one clusters the data in CV space and uses cluster centroids as discrete states of a Markov chain, whose transition rates are estimated at a given lag time. Nevertheless, constructing CVs remains a challenge for real-world macromolecules, although machine learning techniques have been incorporated to develop many useful methods [3, 4]. Therefore, further work is needed to develop a physically grounded, bottom-up stochastic gene expression model.
@@ Appendix A Matrix-form Chemical Master Equation and Generating Function Method
In this section, we show that the chemical master equation (2) can be equivalently reformulated as the partial differential equation system (4) using the standard generating function method.
@@ Appendix D Proof of (14)
In the penultimate inequality, we applied Hölder’s inequality. In the last inequality, we used an analogue to Theorem 4.1.2 in [12], stating that
@@
Since $D$ is the generator of a continuous-time Markov chain, for any $\lambda>0$, the matrix $\lambda\bm{I}_{N}-D$ is strictly row diagonally dominant, with each row summing to $\lambda$. Therefore $\lVert(\lambda\bm{I}_{N}-D)^{-1}\rVert_{\infty}\leq 1/\lambda$.
@@ Appendix E Analytical Expression of Generating Function
where $H(s;z,w;t):=D_{0}+h(s;z,w;t)D_{1}\in\mathbb{R}^{N\times N}$ and $h(s;z,w;t)$ is a real-valued function whose expression is given below.
@@ Appendix F Hierarchical solver for $\mathcal{B}_{p,q}$ by Layers
3:function SolveRight(M,YM,\,Y) $\triangleright$ Return $X$ such that XM=YX\,M=Y (Use LU factorization [12])
@@
12:  $Z_{L}\leftarrow\textsc{SolveRight}(M_{L},\ -\,Y_{L-1}D_{1})$